OpenStudy (anonymous):

if an object is propelled upward from a height of s feet at an initial velocity of v feet per second, then its height h after t seconds is given by the equation h=-16t^2+vt+s, where h is in feet. if the object is propelled from a height of 4 feet with an initial velocity of 64 feet pet second, its height height h given by the equation h=-16t^2+64t+4. after how many seconds is the height 64 feet?

OpenStudy (anonymous):

the equation will be in that condition 64=-16t^2+64t+4. =>16t^2-64t+60=0 => 4t^2-16t+15=0 => (2t-3)(2t+5)=0 so t=3/2 or -5/2 since t<0 is impossible physically, the acceptable solution is t=3/2=1.5 sec.

OpenStudy (anonymous):

thank you!! i had finally figured it out. :)